OpenStudy (anonymous):

Solve the equation. -2p/45 = 7/25

OpenStudy (anonymous):

p=-63/10

OpenStudy (anonymous):

multiply by 45, divide by -2

OpenStudy (anonymous):

p = -45* 7/50 = -6.3